OpenStudy (anonymous):

Rationalize the denominator of 8 over 3 square root of -3?

OpenStudy (saifoo.khan):

dont call that square root. it's a cube root.

OpenStudy (saifoo.khan):

\[\Huge \frac{8}{\sqrt[3]{-3}}\]

OpenStudy (anonymous):

Well excuse me... okay...

OpenStudy (saifoo.khan):

you drew that before.

OpenStudy (anonymous):

does it look like \[\frac{8}{3\sqrt{-3}}\] or \[\frac{8}{\sqrt[3]{-3}}\] ????

OpenStudy (anonymous):

Okay, and........ if you aren't going to help then don't respond. Take your negativity and unwanted comments somewhere else. Thanks ahead of time :)

OpenStudy (anonymous):

The first one Tyler is what it looks like :)

OpenStudy (anonymous):

\[\frac{8i\sqrt{3}}{-3}\]

OpenStudy (anonymous):

I think thats right.

OpenStudy (anonymous):

yeah, but the only choices I have that are given as answers are as follows

OpenStudy (saifoo.khan):

lol @ negativity comments. :P

OpenStudy (anonymous):